Systems Staff Meeting Notes 1/10/00

Systems Staff Meeting Notes
January 10, 2000

Present: Rich Bennett, Will Chaney, Bill Covey, Debra Harris, Winston Harris, Bobbie Meng, Suzy Shaw, Hilton Soto, Ying Tang
Announcements
  1. Pest Control has been sent for (again) - the varmint tried to eat through the refrigerator to get to the doughnuts.
  2. A question was posed about proprietary NOTIS database access for courtesy patrons - whether it would be possible for them to buy access. The answer is no, we are unable to legally sell access to a database that we license.
  3. Bobbie received an email from Everett, he did well in his race in San Diego and said the Pacific Ocean is cold!
  4. Hilton got a new girlfriend over the weekend.
  5. Rich got a new roof on his house.
Review Minutes of 1/3/00
  1. Networking #11 - Bill has not received any feedback regarding space requirements for server applications.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Orders: space in the vault has greatly improved - there are only 4 out of 60 Cows left. There hasn't been much time to receive the old machines & room 2 is full. Bobbie is trying to work out an arrangement with Bill Hanssen whereby they will both meet where the machines to be picked up are located, check off the machines that should be surveyed and then Bill will deliver them to Property.
  2. Price Wars - good prices will facilitate the mad pace of upgrades. Expect that we will continue to purchase 60 Cows at a time, replacing 75's & 100's. Inside deliveries are not an option, our orders will get unloaded by Facilities.
Networking
  1. Need to check on Saturday's wiring project in AFA.
  2. Network backups - Accelerator agent has been installed on NT4, NT5, NT6 & NT8. Backup is now running during the day, will move to nighttime.
  3. Coordination of network wiring projects/phones with Facilities - Facilities will be approached first, they will check with Systems then proceed from there. Bill is writing up this proposal.
  4. SNA timeouts - one person has said the change to the switch has made no difference. Bill believes it may be a FCLA/NERDC problem.
  5. West wiring - on hold until the building construction has been decided.
  6. Norman Hall - wiring project has not started yet. Music, Norman and 203a East are all on the same PO.
Software Projects Status
  1. Student-Patron Updates have been completed.
  2. FCLA backups - there have been some problems, the first tape has not been created yet. Winston has checked with Pat Crabb, it may not be done until this coming weekend. FCLA had a higher number of general failures last week.
  3. Wiring database - still waiting on Bobbie's feedback.
  4. Trouble Ticket Viewer - need to announce it's availability to the Liaisons & post the URL on the Liaison page on the Systems homepage.
  5. OCLC tape conversion is finished, Gerald is inputting through ONI.
  6. Preservation's Q&A database should be moved up to Current Projects.
  7. PAULA application on NT 7 needs to be moved to NT4.
Calendar
  1. MLK Holiday is Jan. 17.
  2. Debra out Jan. 20-27.
  3. Rich out to ALA in San Antonio Jan. 14-19
  4. Everett out Jan. 28-31.


Back to the Systems Homepage

Last updated January 31, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu